A simple TEX Example

Alice was beginning to get very tired of
sitting by her sister on the
bank, and of having nothing to do: once
or twice she had peeped into the book her
sister was reading, but it had no pictures
or conversations in it, `and what is the
use of a book,' thought Alice
`without pictures or conversation?'
\end

Alice was beginning to get very tired of sitting by her sister on the bank, and of having nothing to do: once or twice she had peeped into the book her sister was reading, but it had no pictures or conversations in it, ‘and what is the use of a book,’ thought Alice ‘without pictures or conversation?’